The left and right page numbers of an open book are two consecutive integers whose sum is 419. Find the page numbers
1
Expert's answer
2020-09-24T17:06:09-0400

Left page = x

Right page = x + 1

"x + (x + 1) = 419"

"2*x + 1 = 419"

"2*x = 418"

"x = 418\/2 = 209"

Answer:

Left page = 209,

Right page = 210.